OpenStudy (anonymous):

I am trying to evaluate 3/x+2=2/x-6 do I set these to 0 and solve ?

OpenStudy (anonymous):

you can but it easier to combine like terms on either side of the equal sign

OpenStudy (anonymous):

first you can cross-multiply

OpenStudy (anonymous):

so 2(x+2) = 3(x - 6) then its easy

OpenStudy (anonymous):

expand the brackets and bring like terms either side of the =

OpenStudy (anonymous):

I crossed multiply and did not get what I thought it came out like 2=3 and I cannot find it in my book! but as I was looking I was asked to solve for F=KMn over R^2 what is that suppose to find? My previous problem was 3 over x+2= 2 over x-6